OpenStudy (anonymous):

find the equation Q=P(t)=ab^t, that begins at 6000and quadruple every 2 years

OpenStudy (dumbcow):

"begins at 6,000" is the initial value when t=0 P(0) = a = 6000 so a = 6000 quadruples every 2 years so P(2) = 4*P(0) --> 6000b^2 = 4*6000 solve for b